Use the discriminant to determine whether the following equations have solutions that are: two different rational solutions; two different irrational solutions; exactly one rational solution; or two different imaginary solutions.
x2 = -4x – 4
Two different rational solutions
Exactly one rational solution
Two different irrational solutions
Two different imaginary solutions

Answer by stanbon(75887)

Use the discriminant to determine whether the following equations have solutions that are: two different rational solutions; two different irrational solutions; exactly one rational solution; or two different imaginary solutions.
x^2 = -4x – 4
---
Rearrange:
x^2+4x+4 = 0
a = 1 ; b = 4 ; c = 4
---
b^2 -4ac = 16-4*1*4 = 0
---
Since the discriminant is zero
there is exactly one rational solution.
======================
Cheers,
stan H.
======================
Two different rational solutions
Exactly one rational solution
Two different irrational solutions
Two different imaginary solutions